# Spectral Representation-based Reinforcement Learning

### v1

@@ 4 Algorithms for Learning Spectral Representations
We begin by estimating the representation $\bm{\phi}_{\text{Linear}}$ in the context of linear formulation (11). Several works on linear MDPs (agarwal2020flambe; uehara2021representation) derive the representation by solving the following maximum log-likelihood estimation (MLE) problem:
@@ 4.1 Representations from Spectral Contrastive Learning
Apparently, the triple integral over spaces makes it an intractable objective. Instead, Speder (ren2023spectral) considers a parameterization modulated by a scaling factor $\mathbb{P}(s^{\prime})$:
@@
As shown later, as long as $\mathbb{P}(s^{\prime})$ is amenable to sampling, introducing this factor leads to a tractable optimization objective without altering the final representation. To train $\bm{\varphi}_{\theta}$ and $\bm{\nu}_{\theta}$, Speder leverages spectral contrastive learning (haochen2021provable; ren2023spectral) to match a rebalanced version of (26):
@@ 4.2 Representations from Variational Learning
When the transition operator possesses the latent variable structure as described in Section 3.2.2, LV-Rep (ren2023latent) employs variational learning to obtain a tractable surrogate objective of the original MLE (24). Specifically, suppose the transition operator is parameterized as $\mathbb{P}(s^{\prime}|s,a)=\int\varphi_{\theta}(z|s,a)\nu_{\theta}(s^{\prime}|z)\mathrm{d}z$ for some valid probability measure $\varphi_{\theta}:{\mathcal{S}}\times\mathcal{A}\to\Delta(\mathcal{Z})$, then it follows from Jensen’s inequality that
@@
where $q_{\theta}(\cdot|s,a,s^{\prime})$ is the variational distribution also parameterized by $\theta$, and the RHS of the last inequality is also known as the evidence lower bound (ELBO) (kingma2013auto). We can also introduce a hyper-parameter $\beta\geqslant 0$ and optimize the $\beta$-VAE objective (higgins2017beta) to balance the KL regularization term and the reconstruction term:
@@ 4.3 Representations from Score Matching
In section 3.2.3, we described how to construct spectral representations $\bm{\phi}_{\text{EBM}}$ from a factorized energy function $\bm{\varphi}(s,a)^{\top}\bm{\nu}(s^{\prime})$. A key challenge with this is that estimating the energy function of EBMs with MLE requires computing the normalization factor, which is often intractable (lecun2006tutorial; song2021train). Diffusion models (ho2020denoising; songscore) offer a promising solution for sampling from EBMs by estimating the score functions instead of the energy functions, thereby bypassing the need for this normalization factor. Inspired by this, Diff-SR (shribak2024diffusion) derives a tractable, score-based optimization objective for learning $\bm{\phi}(s,a)$.
@@
which shares the same optimal solution to $\ell_{\mathrm{SM}}$, as proved in the Appendix A of shribak2024diffusion. Note that the regression target has analytical forms when the perturbation kernel is known, and in our case, $\nabla_{\tilde{s}^{\prime}}\log\mathbb{P}(\tilde{s}^{\prime}|s^{\prime};\beta)=-(\tilde{s}^{\prime}-\sqrt{1-\beta}s^{\prime})/{\beta}$. In practice, Diff-SR directly parameterizes $\nabla_{\tilde{s}^{\prime}}\bm{\nu}_{\theta}(\tilde{s}^{\prime};\beta)$ as a neural network $\bm{\kappa}_{\theta}:\mathbb{R}^{\dim({\mathcal{S}})}\times\mathbb{R}\to\mathbb{R}^{d\times\dim({\mathcal{S}})}$ and thereby avoids the need to compute second-order gradients.
@@ 4.4 Representations from Noise Contrastive Estimation
Another effective approach to learning the energy function is noise contrastive estimation (NCE) (ma2018noise; gutmann2010noise; gutmann2012noise). Specifically, let $\widetilde{\mathbb{P}}(s^{\prime})$ be an arbitrary noise distribution, and we model the target density as:
@@
However, when the negative sample distribution $\widetilde{\mathbb{P}}(s^{\prime})$ is trivial or disjoint from the true conditional density $\mathbb{P}(s^{\prime}|s,a)$, NCE may yield poor representation since classifying between the positives and the negatives is too easy (robinson2020contrastive; rhodes2020telescoping). While existing methods often address this by using more negative samples (radford2021learning; chen2022we) or designing more sophisticated noise distributions (finn2016connection), we propose a simple solution inspired by diffusion models.
@@
CTRL (zhang2022making) leverages the NCE framework to derive linear spectral representations. It formulates the target density as:
@@ 5 Reinforcement Learning with Spectral Representations
Following standard practice, we apply the double Q-network trick (fujimoto2018addressing) to stabilize learning and combat over-estimation bias. Specifically, we maintain two independent set of parameters $\xi_{1},\xi_{2}$ and their exponential moving average (EMA) version $\bar{\theta}_{1},\bar{\theta}_{2}$, and update the $Q$-value functions using the standard TD learning objective:
@@
The optimistic exploration with spectral representation has been justified rigorously in (ren2023latent; ren2023spectral), which we briefly introduce here. We first make the following assumptions about the candidate class and normalization conditions, which are common among similar analysis.
@@
Unlike the linear MDP case (jin2020provably; uehara2021representation), we consider representations in the RKHS $\mathcal{H}_{k}$, which includes all formulations in Section 3 and provides an abstraction. For example, the energy-based formulation can be recognized as specifying $k$ as the Gaussian kernel (17), while the linear formulation can be recovered using a linear kernel.
@@
This theorem and its proof largely follow Theorem 4 in ren2023latent, except that the bonus functions $\hat{b}_{n}$ are defined with the RKHS inner product $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}_{k}}$, rather than $\left\langle\cdot,\cdot\right\rangle_{L_{2}(\mu)}$. We refer interested readers to Appendix E of ren2023latent for a detailed proof.
@@ 6 Spectral Representations in Partially Observable MDPs
In this section, we demonstrate that the framework of spectral representations can be seamlessly extended to partially observable MDPs (POMDPs), leading to the first practical algorithm that can provably solves a subclass of POMDPs. Following efroni2022provable, a POMDP is defined by the tuple $\langle{\mathcal{S}},\mathcal{A},\mathcal{O},\mathbb{P},\mathbb{O},r,\gamma,d_{0}\rangle$, where ${\mathcal{S}},\mathcal{A},$ and $\mathcal{O}$ are the state, action, and observation spaces, $r:\mathcal{O}\times\mathcal{A}\to\mathbb{R}$ is the reward function, $\mathbb{P}:{\mathcal{S}}\times\mathcal{A}\to\Delta({\mathcal{S}})$ is the transition probability over the latent states, and $\mathbb{O}:{\mathcal{S}}\to\Delta(\mathcal{O})$ is the emission probability that governs the observation an agent will receive given the system state.
@@
Definition 6.5.
($L$-decodability (efroni2022provable)) Define
@@
Intuitively, $L$-decodability implies that the $L$-step history window serves as a sufficient statistic for the state $s_{t}$. We note that the $L$-decodability is a plausible assumption that commonly holds in practical decision-making scenarios. For example, in visual RL, although a single frame may be insufficient to capture the full system state (e.g., velocity of the robot), stacking consecutive frames effectively compensates for this partial observability and has become the de facto practice in modern visual RL algorithms (mnih2013playing; yarats2021mastering).
@@
where $V^{\pi}(x_{t+L})$ is independent of $(x_{t},a_{t})$ thanks to the $L$-decodability property. To further eliminate the dependence on $x_{t}$ in $x_{t+1:t+L-1}$, we employ the moment matching policy (efroni2022provable) $\tilde{\pi}$, which conditions solely on $(x_{t},a_{t})$ but generates the same $L$-step observation distribution as $\pi$ (proof provided in zhang2023provable, Appendix C). In this way, if we consider the spectral decomposition of the $L$-step transition and reward:
@@ 7 Empirical Evaluations
In this section, we evaluate the effectiveness of spectral representation-based RL algorithms using the widely recognized DeepMind Control (DMControl) Suite (tassa2018deepmind). Our experiments are designed to answer two primary questions: 1) whether learning spectral representations improves policy optimization, and 2) how different formulations and learning methods for these representations compare against each other and state-of-the-art baselines. To fulfill this, we conduct our evaluation on over 20 distinct tasks from the suite, encompassing both low-dimensional proprioceptive states and high-dimensional visual inputs. Rendered illustrations of these tasks are shown in Figure 3.
@@
    Speder (ren2023spectral), which employs a linear formulation and learns representations $\bm{\phi}_{\text{Linear}}$ via spectral contrastive learning (27);
@@
    LV-Rep (ren2023latent), which trains a latent variable spectral representation $\bm{\phi}_{\text{LV}}$ through variational learning (32);
@@
    Diff-SR (shribak2024diffusion), which utilizes an energy-based formulation $\bm{\phi}_{\text{EBM}}$ optimized with score matching (36);
@@
    CTRL-SR, which improves CTRL (zhang2022making) by using ranking-based perturbed NCE (40) as its representation learning objectives. Note that it also employs the energy-based representations $\bm{\phi}_{\text{EBM}}$ instead of the $\bm{\phi}_{\text{Linear}}$ in CTRL.
@@ 7.1 Results with Proprioceptive Observations
For our baselines, we select two model-free algorithms, TD3 (fujimoto2018addressing) and SAC (haarnoja2018soft). We also include TD7 (fujimoto2023sale), a recent TD3 variant that incorporates representation learning, to provide a strong comparison. A consistent training and evaluation protocol is applied to all algorithms and tasks. Each agent interacts with the environment for 1 million frames, with each action being repeated for two consecutive frames (i.e., an action repeat of two frames). The agent’s policy is updated every two frames, which amounts to 500,000 gradient steps in total. Performance is evaluated every 10,000 frames by averaging the returns over 10 episodes. To ensure statistical reliability, we report the mean and standard deviation of these evaluation scores across 5 independent runs, each initialized with a different random seed.
@@ 7.2 Results with Visual Observations
For the baseline algorithms, we include: 1) DrQ-V2 (yarats2021mastering), a competitive model-free visual RL algorithm which employs data augmentation to enhance the visual encoders; 2) TDMPC2 (hansen2023td), a model-based algorithm that performs model predictive control during sampling; and 3) Dreamer-V3 (hafner2023mastering), the state-of-the-art model-based algorithm that employs an RSSM structured model for agent training.
We implemented the best-performing methods, namely Diff-SR and CTRL-SR, for this setting. To handle the visual input, we adopt a convolusional visual encoder to extract a lower-dimensional latent vector from the stacked frames, and the representation networks treat the extracted latent vectors as states. For CTRL-SR, the visual encoder is jointly optimized along with the representation networks, while for Diff-SR, we found that joint optimization may lead to representation collapse, a similar phenomenon that is also observed by other representation learning methods (grill2020bootstrap; bardes2021vicreg). Thus, we decided to train the visual encoder using a variance-regularized reconstruction objective. More details about the implementation can be found in Appendix B.
@@ 8.1 Reinforcement Learning with Representation Learning
Given the advancements in self-supervised learning (SSL) and the inherent challenges of reinforcement learning (RL) in high-dimensional spaces, there is growing interest in leveraging representation learning to improve RL efficiency. A straightforward application of this approach has been in visual RL, where agents perceive the environment through camera images that often contain redundant information. Many established visual representation techniques, including data augmentation (yarats2021image; yarats2021mastering), observation reconstruction (yarats2021improving), and contrastive representation learning (laskin2020curl), can be readily integrated to derive compact features for the learning.
@@
In parallel to visual representation learning, a long-standing topic in RL is learning temporal representations that encode system behavior and facilitate long-term planning. Among them, successor features (dayan1993improving; kulkarni2016deep; barreto2017successor) and forward-backward representations (touati2021learning) learn to summarize the long-term consequences of actions and can potentially adapt to different reward functions by applying the learned representations. Methods based on bisimulation metrics (ferns2004metrics; gelada2019deepmdp; castro2020scalable; liao2023policy) quantify the behavioral similarity between states and learn representations that group equivalent states together. Inspired by model-based approaches, a large body of work (munk2016learning; lee2020stochastic; guo2020bootstrap; schwarzer2020data; mcinroe2021learning; yu2022mask; tang2023understanding; kim2025self; fujimoto2023sale; fujimoto2025towards; assran2025v) learns representations by predicting future system states, either directly from observations or in a latent space. The spectral representation framework (ren2022free; ren2023latent; zhang2022making; ren2023spectral; shribak2024diffusion) addressed in this paper is closely related to these methods, especially fujimoto2025towards, which learn a linear, self-predictive latent representation of states and actions and build the Q-value function upon them. To see this connection, the second term of the spectral contrastive loss in (27) is equivalent to predicting the representation of the next state under deterministic MDPs and certain normalization conditions on the representations. The variational learning objective in (32) can also be cast as a non-linear extension of the latent representation prediction problem. Finally, the contrastive learning objective used by CTRL-SR has also been employed to extract temporal abstractions of the system by distinguishing ground-truth state transitions from faked ones (oord2018representation; anand2019unsupervised; stooke2021decoupling; zheng2023texttt).
@@
While the learning objectives share many similarities, spectral representations go beyond merely capturing the predictive structure. By exploiting the spectrum structure of the dynamics, spectral representations are distinguished from these alternatives by their unique connection to the $Q$-value functions, which we have demonstrated in Theorem 5.4 is crucial for exploration design and efficient learning.

### v2

@@ 4 Algorithms for Learning Spectral Representations
We begin by estimating the representation $\bm{\phi}_{\text{Linear}}$ in the context of linear formulation (11). Several works on linear MDPs (Agarwal et al., 2020; Uehara et al., 2021) derive the representation by solving the following maximum log-likelihood estimation (MLE) problem:
@@ 4.1 Representations from Spectral Contrastive Learning
Apparently, the triple integral over spaces makes it an intractable objective. Instead, Speder (Ren et al., 2023b) considers a parameterization modulated by a scaling factor $\mathbb{P}(s^{\prime})$:
@@
As shown later, as long as $\mathbb{P}(s^{\prime})$ is amenable to sampling, introducing this factor leads to a tractable optimization objective without altering the final representation. To train $\bm{\varphi}_{\theta}$ and $\bm{\nu}_{\theta}$, Speder leverages spectral contrastive learning (HaoChen et al., 2021; Ren et al., 2023b) to match a rebalanced version of (26):
@@ 4.2 Representations from Variational Learning
When the transition operator possesses the latent variable structure as described in Section 3.2.2, LV-Rep (Ren et al., 2023a) employs variational learning to obtain a tractable surrogate objective of the original MLE (24). Specifically, suppose the transition operator is parameterized as $\mathbb{P}(s^{\prime}|s,a)=\int\varphi_{\theta}(z|s,a)\nu_{\theta}(s^{\prime}|z)\mathrm{d}z$ for some valid probability measure $\varphi_{\theta}:{\mathcal{S}}\times\mathcal{A}\to\Delta(\mathcal{Z})$, then it follows from Jensen’s inequality that
@@
where $q_{\theta}(\cdot|s,a,s^{\prime})$ is the variational distribution also parameterized by $\theta$, and the RHS of the last inequality is also known as the evidence lower bound (ELBO) (Kingma and Welling, 2013). We can also introduce a hyper-parameter $\beta\geqslant 0$ and optimize the $\beta$-VAE objective (Higgins et al., 2017) to balance the KL regularization term and the reconstruction term:
@@ 4.3 Representations from Score Matching
In section 3.2.3, we described how to construct spectral representations $\bm{\phi}_{\text{EBM}}$ from a factorized energy function $\bm{\varphi}(s,a)^{\top}\bm{\nu}(s^{\prime})$. A key challenge with this is that estimating the energy function of EBMs with MLE requires computing the normalization factor, which is often intractable (LeCun et al., 2006; Song and Kingma, 2021). Diffusion models (Ho et al., 2020; Song et al., ) offer a promising solution for sampling from EBMs by estimating the score functions instead of the energy functions, thereby bypassing the need for this normalization factor. Inspired by this, Diff-SR (Shribak et al., 2024) derives a tractable, score-based optimization objective for learning $\bm{\phi}(s,a)$.
@@
which shares the same optimal solution to $\ell_{\mathrm{SM}}$, as proved in the Appendix A of Shribak et al. (2024). Note that the regression target has analytical forms when the perturbation kernel is known, and in our case, $\nabla_{\tilde{s}^{\prime}}\log\mathbb{P}(\tilde{s}^{\prime}|s^{\prime};\beta)=-(\tilde{s}^{\prime}-\sqrt{1-\beta}s^{\prime})/{\beta}$. In practice, Diff-SR directly parameterizes $\nabla_{\tilde{s}^{\prime}}\bm{\nu}_{\theta}(\tilde{s}^{\prime};\beta)$ as a neural network $\bm{\kappa}_{\theta}:\mathbb{R}^{\dim({\mathcal{S}})}\times\mathbb{R}\to\mathbb{R}^{d\times\dim({\mathcal{S}})}$ and thereby avoids the need to compute second-order gradients.
@@ 4.4 Representations from Noise Contrastive Estimation
Another effective approach to learning the energy function is noise contrastive estimation (NCE) (Ma and Collins, 2018; Gutmann and Hyvärinen, 2010, 2012). Specifically, let $\widetilde{\mathbb{P}}(s^{\prime})$ be an arbitrary noise distribution, and we model the target density as:
@@
However, when the negative sample distribution $\widetilde{\mathbb{P}}(s^{\prime})$ is trivial or disjoint from the true conditional density $\mathbb{P}(s^{\prime}|s,a)$, NCE may yield poor representation since classifying between the positives and the negatives is too easy (Robinson et al., 2020; Rhodes et al., 2020). While existing methods often address this by using more negative samples (Radford et al., 2021; Chen et al., 2022b) or designing more sophisticated noise distributions (Finn et al., 2016), we propose a simple solution inspired by diffusion models.
@@
CTRL (Zhang et al., 2022) leverages the NCE framework to derive linear spectral representations. It formulates the target density as:
@@ 5 Reinforcement Learning with Spectral Representations
Following standard practice, we apply the double Q-network trick (Fujimoto et al., 2018) to stabilize learning and combat over-estimation bias. Specifically, we maintain two independent set of parameters $\xi_{1},\xi_{2}$ and their exponential moving average (EMA) version $\bar{\theta}_{1},\bar{\theta}_{2}$, and update the $Q$-value functions using the standard TD learning objective:
@@
The optimistic exploration with spectral representation has been justified rigorously in (Ren et al., 2023a, b), which we briefly introduce here. We first make the following assumptions about the candidate class and normalization conditions, which are common among similar analysis.
@@
Unlike the linear MDP case (Jin et al., 2020; Uehara et al., 2021), we consider representations in the RKHS $\mathcal{H}_{k}$, which includes all formulations in Section 3 and provides an abstraction. For example, the energy-based formulation can be recognized as specifying $k$ as the Gaussian kernel (17), while the linear formulation can be recovered using a linear kernel.
@@
This theorem and its proof largely follow Theorem 4 in Ren et al. (2023a), except that the bonus functions $\hat{b}_{n}$ are defined with the RKHS inner product $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}_{k}}$, rather than $\left\langle\cdot,\cdot\right\rangle_{L_{2}(\mu)}$. We refer interested readers to Appendix E of Ren et al. (2023a) for a detailed proof.
@@ 6 Spectral Representations in Partially Observable MDPs
In this section, we demonstrate that the framework of spectral representations can be seamlessly extended to partially observable MDPs (POMDPs), leading to the first practical algorithm that can provably solves a subclass of POMDPs. Following Efroni et al. (2022), a POMDP is defined by the tuple $\langle{\mathcal{S}},\mathcal{A},\mathcal{O},\mathbb{P},\mathbb{O},r,\gamma,d_{0}\rangle$, where ${\mathcal{S}},\mathcal{A},$ and $\mathcal{O}$ are the state, action, and observation spaces, $r:\mathcal{O}\times\mathcal{A}\to\mathbb{R}$ is the reward function, $\mathbb{P}:{\mathcal{S}}\times\mathcal{A}\to\Delta({\mathcal{S}})$ is the transition probability over the latent states, and $\mathbb{O}:{\mathcal{S}}\to\Delta(\mathcal{O})$ is the emission probability that governs the observation an agent will receive given the system state.
@@
Definition 4
($L$-decodability (Efroni et al., 2022)) Define
@@
Intuitively, $L$-decodability implies that the $L$-step history window serves as a sufficient statistic for the state $s_{t}$. We note that the $L$-decodability is a plausible assumption that commonly holds in practical decision-making scenarios. For example, in visual RL, although a single frame may be insufficient to capture the full system state (e.g., velocity of the robot), stacking consecutive frames effectively compensates for this partial observability and has become the de facto practice in modern visual RL algorithms (Mnih et al., 2013; Yarats et al., 2021a).
@@
where $V^{\pi}(x_{t+L})$ is independent of $(x_{t},a_{t})$ thanks to the $L$-decodability property. To further eliminate the dependence on $x_{t}$ in $x_{t+1:t+L-1}$, we employ the moment matching policy (Efroni et al., 2022) $\tilde{\pi}$, which conditions solely on $(x_{t},a_{t})$ but generates the same $L$-step observation distribution as $\pi$ (proof provided in Zhang et al. , Appendix C). In this way, if we consider the spectral decomposition of the $L$-step transition and reward:
@@ 7 Empirical Evaluations
In this section, we evaluate the effectiveness of spectral representation-based RL algorithms using the widely recognized DeepMind Control (DMControl) Suite (Tassa et al., 2018). Our experiments are designed to answer two primary questions: 1) whether learning spectral representations improves policy optimization, and 2) how different formulations and learning methods for these representations compare against each other and state-of-the-art baselines. To fulfill this, we conduct our evaluation on over 20 distinct tasks from the suite, encompassing both low-dimensional proprioceptive states and high-dimensional visual inputs. Rendered illustrations of these tasks are shown in Figure 3.
@@
    Speder (Ren et al., 2023b), which employs a linear formulation and learns representations $\bm{\phi}_{\text{Linear}}$ via spectral contrastive learning (27);
@@
    LV-Rep (Ren et al., 2023a), which trains a latent variable spectral representation $\bm{\phi}_{\text{LV}}$ through variational learning (32);
@@
    Diff-SR (Shribak et al., 2024), which utilizes an energy-based formulation $\bm{\phi}_{\text{EBM}}$ optimized with score matching (36);
@@
    CTRL-SR, which improves CTRL (Zhang et al., 2022) by using ranking-based perturbed NCE (40) as its representation learning objectives. Note that it also employs the energy-based representations $\bm{\phi}_{\text{EBM}}$ instead of the $\bm{\phi}_{\text{Linear}}$ in CTRL.
@@ 7.1 Results with Proprioceptive Observations
For our baselines, we select two model-free algorithms, TD3 (Fujimoto et al., 2018) and SAC (Haarnoja et al., 2018). We also include TD7 (Fujimoto et al., 2023), a recent TD3 variant that incorporates representation learning, to provide a strong comparison. A consistent training and evaluation protocol is applied to all algorithms and tasks. Each agent interacts with the environment for 1 million frames, with each action being repeated for two consecutive frames (i.e., an action repeat of two frames). The agent’s policy is updated every two frames, which amounts to 500,000 gradient steps in total. Performance is evaluated every 10,000 frames by averaging the returns over 10 episodes. To ensure statistical reliability, we report the mean and standard deviation of these evaluation scores across 5 independent runs, each initialized with a different random seed.
@@ 7.2 Results with Visual Observations
For the baseline algorithms, we include: 1) DrQ-V2 (Yarats et al., 2021a), a competitive model-free visual RL algorithm which employs data augmentation to enhance the visual encoders; 2) TDMPC2 (Hansen et al., 2023), a model-based algorithm that performs model predictive control during sampling; and 3) Dreamer-V3 (Hafner et al., 2023), the state-of-the-art model-based algorithm that employs an RSSM structured model for agent training.
We implemented the best-performing methods, namely Diff-SR and CTRL-SR, for this setting. To handle the visual input, we adopt a convolusional visual encoder to extract a lower-dimensional latent vector from the stacked frames, and the representation networks treat the extracted latent vectors as states. For CTRL-SR, the visual encoder is jointly optimized along with the representation networks, while for Diff-SR, we found that joint optimization may lead to representation collapse, a similar phenomenon that is also observed by other representation learning methods (Grill et al., 2020; Bardes et al., 2021). Thus, we decided to train the visual encoder using a variance-regularized reconstruction objective. More details about the implementation can be found in Appendix B.
@@ 8.1 Reinforcement Learning with Representation Learning
Given the advancements in self-supervised learning (SSL) and the inherent challenges of reinforcement learning (RL) in high-dimensional spaces, there is growing interest in leveraging representation learning to improve RL efficiency. A straightforward application of this approach has been in visual RL, where agents perceive the environment through camera images that often contain redundant information. Many established visual representation techniques, including data augmentation (Yarats et al., 2021b, a), observation reconstruction (Yarats et al., 2021c), and contrastive representation learning (Laskin et al., 2020), can be readily integrated to derive compact features for the learning.
@@
In parallel to visual representation learning, a long-standing topic in RL is learning temporal representations that encode system behavior and facilitate long-term planning. Among them, successor features (Dayan, 1993; Kulkarni et al., 2016; Barreto et al., 2017) and forward-backward representations (Touati and Ollivier, 2021) learn to summarize the long-term consequences of actions and can potentially adapt to different reward functions by applying the learned representations. Methods based on bisimulation metrics (Ferns et al., 2004; Gelada et al., 2019; Castro, 2020; Liao et al., 2023) quantify the behavioral similarity between states and learn representations that group equivalent states together. Inspired by model-based approaches, a large body of work (Munk et al., 2016; Lee et al., 2020; Guo et al., 2020; Schwarzer et al., 2020; McInroe et al., 2021; Yu et al., 2022; Tang et al., 2023; Kim et al., 2025; Fujimoto et al., 2023, 2025; Assran et al., 2025) learns representations by predicting future system states, either directly from observations or in a latent space. The spectral representation framework (Ren et al., 2022, 2023a; Zhang et al., 2022; Ren et al., 2023b; Shribak et al., 2024) addressed in this paper is closely related to these methods, especially Fujimoto et al. (2025), which learn a linear, self-predictive latent representation of states and actions and build the Q-value function upon them. To see this connection, the second term of the spectral contrastive loss in (27) is equivalent to predicting the representation of the next state under deterministic MDPs and certain normalization conditions on the representations. The variational learning objective in (32) can also be cast as a non-linear extension of the latent representation prediction problem. Finally, the contrastive learning objective used by CTRL-SR has also been employed to extract temporal abstractions of the system by distinguishing ground-truth state transitions from faked ones (Oord et al., 2018; Anand et al., 2019; Stooke et al., 2021; Zheng et al., 2023).
@@
While the learning objectives share many similarities, spectral representations go beyond merely capturing the predictive structure. By exploiting the spectrum structure of the dynamics, spectral representations are distinguished from these alternatives by their unique connection to the $Q$-value functions, which we have demonstrated in Theorem 3 is crucial for exploration design and efficient learning.